OpenStudy (anonymous):

12 less than twice the square of a number is equal to the product of -5 and the number. What are the solutions? a) {-3/2 , -4} b) { -3/2, 4} c) {3/2, 4} d) {3/2, -4}

OpenStudy (anonymous):

d \2[x^2-12=-5x\] \[2x^2+5x-12=0\]

OpenStudy (anonymous):

\[2x^2-12=-5x\] \[2x^2+5x-12=0\] \[(x+4)(2x-3)=0\] etc

OpenStudy (anonymous):

Thank you so much! I really suck at math and i dont know where i would be without a website like this. THANK YOU! (: